Ignoring leap year, if the probability of anyone person’s birthday is 1/365, then the probability of another having a different birthday would be 1/364, and a third having a unique birthday would be 1/363, etc. So in a group of 30, what would be the probability of any two people having the same birthday?
Answer #1

probability of any two people having the same birthday =1-P(all 30 have different birthdays)

=1-(365*364*363*362.........(365-30+1))/36530 =1-0.293684 =0.706316